Prove that





gamma=(a(3u,u^ 8 ), 3a * u ^ 2, 3a * (3u + u ^ 3)





tau = 1/(k ^ 2) * (r*gamma)

1
Expert's answer
2021-11-15T17:13:33-0500

"\gamma=(a(3u+u^ 8 ), 3a u ^ 2, 3a (3u + u ^ 3))"


"\tau=\frac{r\gamma}{k^2}"


torsion:


"\tau=\frac{(\gamma'\times \gamma'')\cdot\gamma'''}{|\gamma'\times \gamma''|^2}"


curvature:


"k=\frac{|\gamma'\times \gamma''|}{|\gamma'|^3}"


then:

"(\gamma'\times \gamma'')\cdot\gamma'''=|\gamma'|^6r\gamma"



"\gamma'=(a(3+8u^ 7 ), 6au , 3a (3 + 3u ^2))"

"\gamma''=(56au^ 6 , 6a , 18au)"

"\gamma'''=(336au^ 5 , 0 , 18a)"


"(\gamma'\times \gamma'')\cdot\gamma'''=\begin{vmatrix}
 a(3+8u^ 7 ) & 6au& 3a (3 + 3u ^2)\\
 56au^ 6 & 6a&18au\\
336au^ 5&0&18a
\end{vmatrix}="


"=-6au(1008a^2u^6-6048a^2u^6)+6a(54a^2+144a^2u^7-3024a^2u^5-3024a^2u^7)="

"=12960a^3u^7-18144a^3u^5+324a^3"


"|\gamma'|^6=((3a+8au^ 7 )^2+36a^2u ^ 2+(9a + 9au ^ 2))^3"